# Countably compact inverse semigroups and Nyikos problem

### v1

@@ 2. Compact-like semigroups with dense inverse subsemigroups and the automatic continuity of inversion
$xx^{-1}=xx^{-2}x^{2}x^{-2}x^{2}x^{-1}=xx^{-2}x^{2}x^{-1}=(xx^{-1})(x^{-1}x)(xx^{-1})=(xx^{-1})(x^{-1}x).$
Hence $xx^{-1}=x^{-1}x$ and, as such, the semigroup $S$ is Clifford. ∎
By $\beta X$ we denote the Stone-Čech compactification of a Tychonoff space $X$ (see [22, Chapter 3.6]). The following result appears in [3, Theorem 2.3].
@@
We are in a position to prove Theorem D. We need to show that inversion is continuous in any inverse topological semigroup $S$ provided (i) $S{\times}S$ is pseudocompact, or (ii) $S$ is regular, topologically periodic and $S{\times}S$ is countably compact.
Proof of Theorem D.
(i) By Theorem 2.9, $\beta S$ is a compact topological semigroup that contains $S$ as a dense inverse subsemigroup. Theorem C implies that $\beta S$ is a topological inverse semigroup. Then the continuity of inversion in $\beta S$ yields the continuity of inversion in $S$.
@@
3. Chains in topological semilattices
Recall that each semilattice possesses the natural partial order $\leq$ defined by $a\leq b$ if and only if $ab=ba=a$. Further, if a semilattice is treated as a poset, then by default it is assumed to carry the natural partial order. For an element $x$ of a semilattice $X$ put ${\downarrow}x=\{y\in X\colon y\leq x\}$ and ${\uparrow}x=\{y\in X\colon x\leq y\}$. Note that if $X$ is a semitopological semilattice, then for every $x\in X$ the sets ${\downarrow}x$ and ${\uparrow}x$ are closed. A semitopological semilattice $X$ is called chain-compact if all maximal chains in $X$ are compact. The following two theorems follow from [1, Theorem 3.1].
@@ 3. Chains in topological semilattices
Let $X$ be a semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\min)$. If $\overline{L}$ is compact, then there exists $z=\sup L\in\overline{L}$ such that for each open neighborhood $U$ of $z$ the set $L\setminus U$ is countable.
@@
By Proposition 3.3, $\overline{L}$ is a compact chain. Theorem 3.1(2) applied to $\overline{L}$ implies that $\overline{L}$ contains the supremum of $L$ which we denote by $z$. It is clear that $z=\sup L^{\prime}$ for each uncountable subset $L^{\prime}$ of $L$. By Theorem 3.1(2), $z\in\overline{L^{\prime}}$ for each uncountable subset $L^{\prime}$ of $L$. It follows that for each open neighborhood $U$ of $z$ the set $L\setminus U$ is countable, as required. ∎
@@
Let $X$ be a semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\max)$. If $\overline{L}$ is compact, then there exists $z=\inf L\in\overline{L}$ such that for each open neighborhood $U$ of $z$ the set $L\setminus U$ is countable.
@@
Let $L$ be a chain isomorphic to $(\omega_{1},\min)$. If $\overline{L}$ is compact, then Lemma 3.4 yields $z=\sup L\in\overline{L}$. The countable tightness of $X$ implies that there exists a countable subset $A$ of $L$ such that $z\in\overline{A}$. Then there exists $l\in L$ such that $A\subseteq{\downarrow}l$. Since $X$ is a semitopological semilattice, $X\setminus{\downarrow}l$ is an open neighborhood of $z$ disjoint with $A$, a contradiction. In case when $L$ is isomorphic to $(\omega_{1},\max)$, the proof is similar. ∎
We are in a position to prove Theorem B. We need to show that for a Tychonoff Nyikos topological semilattice $X$ the following conditions are equivalent: (i) $X$ is compact; (ii) no chain in $X$ is isomorphic to $(\omega_{1},\min)$ or $(\omega_{1},\max)$; (iii) no chain in $X$ is topologically isomorphic to $(\omega_{1},\min)$ or $(\omega_{1},\max)$ equipped with the order topology.
Proof of Theorem B.
@@
(iii) $\Rightarrow$ (i): Since $X$ is first-countable, for any cardinal $\kappa>\omega_{1}$ there is no chain in $X$ that is topologically isomorphic to $(\kappa,\min)$ or $(\kappa,\max)$ endowed with the order topology. Since $X$ is countably compact, no chain in $X$ is topologically isomorphic to the discrete semilattices $(\omega,\min)$ and $(\omega,\max)$. By the assumption, no chain in $X$ is topologically isomorphic to $(\omega_{1},\min)$ or $(\omega_{1},\max)$ endowed with the order topology. Theorem 3.1(3) implies that the semilattice $X$ is chain-compact. By [22, Corollary 3.10.15], $X{\times}X$ is countable compact and thus pseudocompact. By Theorem 2.9, $\beta X$ is a topological semigroup that contains $X$ as a dense subsemilattice. Theorem 3.2 implies that $X$ is closed in $\beta X$ and thus $X=\beta X$ is compact. ∎
The rest of this section contains technical results on chains in topological semilattices. They will be used to prove the compactness of locally compact Nyikos topological semilattices (see Theorem 4.12), which is a milestone in the proof of Theorem A.
@@
    For each open neighborhood $U$ of $x$ the set $C\setminus U$ is finite;
@@
(2) $\Rightarrow$ (3): Assume that $x\in\overline{C}$. Fix any $c\in C$ and open neighborhood $U$ of $x$. Note that the set ${\uparrow}c\cap U$ is nonempty, witnessing that $c\in cU$. Since $U$ was arbitrarily chosen and $X$ is a Hausdorff semitopological semilattice, we get that $cx=c$. Since the point $c$ was chosen arbitrarily, we obtain that $c\leq x$ for all $c\in C$, i.e. $x$ is an upper bound of $C$. Assume that $y$ is another upper bound of $C$. Taking into account that $x\in\overline{C}$, we have $U\cap yU\neq\varnothing$ for each open neighborhood $U$ of $x$. Since $X$ is a Hausdorff semitopological semilattice, $x=yx$. Hence $x=\sup C$.
(3) $\Rightarrow$ (1): To derive a contradiction, assume that there exists an open neighborhood $U$ of $x$ such that the set $D=C\setminus U$ is infinite. It is clear that $D$ is a cofinal subset of $C$ isomorphic to $(\omega,\min)$. In particular, we have $\sup D=x$. By the countable compactness of $X$, there exists $y\in\overline{D}$. Repeating the arguments above, we obtain $y=\sup D$. Hence $y=x$, but $x\notin\overline{D}$, a contradiction. ∎
@@
Let $X$ be a semitopological semilattice and $L\subset X$ be a chain isomorphic to $(\alpha,\min)$ for some ordinal $\alpha$. If $\overline{L}$ does not contain maximum, then $L$ is cofinal in $\overline{L}$ and $\overline{L}$ is isomorphic to $(\alpha,\min)$.
@@
Let $L=\{l_{\xi}:\xi\in\alpha\}$, where $l_{\xi}\leq l_{\eta}$ if and only if $\xi\leq\eta$. Fix any $x\in\overline{L}$. If $L\subseteq{\downarrow}x$, then $\overline{L}\subseteq{\downarrow}x$, as ${\downarrow}x$ is closed in $X$. But then $x$ is the maximum of $\overline{L}$, a contradiction. So, for each $x\in\overline{L}$, the set $L\setminus{\downarrow}x$ is nonempty. It follows that the ordinal $\alpha$ is limit and $L$ is cofinal in $\overline{L}$.
For every $x\in\overline{L}\setminus L$ let $\delta(x)=\min\{\xi\in\alpha:x\leq l_{\xi}\}$. Since $L$ is cofinal in $\overline{L}$, the ordinals $\delta(x)$, $x\in\overline{L}\setminus L$, are well defined. By Proposition 3.3, $\overline{L}$ is a chain. Seeking a contradiction, assume that there exists an infinite decreasing chain $\{x_{n}:n\in\omega\}\subset\overline{L}$. The chain $L$, being isomorphic to an ordinal, contains only finite decreasing chains. Hence we lose no generality assuming that $\{x_{n}:n\in\omega\}\subset\overline{L}\setminus L$. Observe that $\{\delta(x_{n}):n\in\omega\}$ is a nonincreasing chain in $\alpha$. Thus there exists $k\in\omega$ such that $\delta(x_{n})=\delta(x_{m})$ for each $n,m\geq k$. Then the set $W=X\setminus({\uparrow}x_{k}\cup{\downarrow}x_{k+2})$ is an open neighborhood of $x_{k+1}$ which is disjoint with $L$. The obtained contradiction implies that $\overline{L}$ contains only finite decreasing chains. Hence $\overline{L}$ is isomorphic to $(\theta,\min)$ for some ordinal $\theta$. Fix the order isomorphism $\psi:\overline{L}\rightarrow\theta$. It is easy to see that $\psi$ can be defined recursively as follows: $\psi(l_{0})=0$ and for each $y\in\overline{L}$,
$\psi(y)=\begin{cases}\psi(x)+1,&\hbox{if there exists }x<y\hbox{ such that }\{z\in\overline{L}:x<z<y\}=\varnothing;\\ \sup\{\psi(x):x<y\},&\hbox{otherwise}.\end{cases}$
@@
Let $X$ be a countably compact countably tight semitopological semilattice and $L\subset X$ be a chain isomorphic to $(\omega_{1},\min)$. Then $\overline{L}$ is topologically isomorphic to $(\omega_{1},\min)$ endowed with the order topology, and $L$ is cofinal in $\overline{L}$.
@@
Fix any $x\in\overline{L}\setminus L$. Since $X$ is countably tight, there exists a countable subset $A$ of $L$ such that $x\in\overline{L}$. It is clear that there exists $l\in L$ such that $A\subseteq{\downarrow}l$. Since $X$ is a semitopological semillatice, the set ${\downarrow}l$ is closed and, as such, $x\in\overline{A}\subseteq{\downarrow}l$. It follows that for each $x\in\overline{L}$ there exists $l\in L$ with $x\leq l$. Thus, $\overline{L}$ does not contain maximum. Lemma 3.9 implies that $\overline{L}$ is isomorphic to $(\omega_{1},\min)$ and $L$ is cofinal in $\overline{L}$. Let $\overline{L}=\{y_{\xi}:\xi<\omega_{1}\}$, where $y_{\xi}\leq y_{\delta}$ if and only if $\xi\leq\delta$. Fix any $\xi<\omega_{1}$. Clearly, for each $\delta<\xi$ the set
@@
is open in $\overline{L}$. Hence the order topology on $\overline{L}$ is contained in the original one. To derive a contradiction, assume that the order topology on $\overline{L}$ is strictly coarser than the original one. Then, taking into account that for each successor ordinal $\theta<\omega_{1}$ the point $y_{\theta}$ is isolated in both topologies, there exists a limit ordinal $\xi<\omega_{1}$ and an open neighborhood $U$ of $y_{\xi}$ such that for any $\delta<\xi$ the set $\{y_{\alpha}:\delta<\alpha<\xi+1\}\setminus U$ is not empty. Since $\xi$ is countable, there exists a cofinal in ${\downarrow}y_{\xi}\setminus\{y_{\xi}\}$ subset $\{\pi_{n}:n\in\omega\}\subset\overline{L}\setminus U$. But this contradicts Lemma 3.8, as $y_{\xi}=\sup\{\pi_{n}:n\in\omega\}$, but $y_{\xi}\notin\overline{\{\pi_{n}:n\in\omega\}}$. ∎
4. Nyikos semilattices
@@
Let $D$ be a countable dense subset of a countably compact semitopological semilattice $X$. Recall that for each $x\in X$ the set ${\downarrow}x$ is closed. Since $X$ is countably compact and the family $\{{\downarrow}x:x\in X\}$ is centered, $Z=\bigcap_{d\in D}{\downarrow}d\neq\varnothing$. Pick any $z\in Z$ and note that $zX=z\overline{D}\subset\overline{zD}=\overline{\{z\}}=\{z\}$. Hence $z$ is the minimum of $X$. ∎
@@
For every semilattice $X$ and chain $L\subseteq X$ the set $I_{L}$ is an ideal.
@@
If $I_{L}=\varnothing$, then there is nothing to prove. Fix any $x\in X$ and $y\in I_{L}$. Since $|yL|<|L|$, we get $|xyL|=|x(yL)|\leq|yL|<|L|$. Hence $xy\in I_{L}$, witnessing that $I_{L}$ is an ideal. ∎
@@
By Lemma 4.3, $I_{L}$ is an ideal. Let $L=\{l_{\xi}:\xi<\omega_{1}\}$, where $l_{\xi}\leq l_{\beta}$ if and only if $\xi\leq\beta$. Fix any accumulation point $z$ of $I_{L}$. Since $X$ is countably tight, there exists a countable subset $B\subseteq I_{L}$ such that $z\in\overline{B}$. For each $b\in B$ there exists an ordinal $\alpha_{b}<\omega_{1}$ such that $bl_{\xi}=bl_{\alpha_{b}}$ for all $\xi\geq\alpha_{b}$. Let $\alpha=\sup\{\alpha_{b}:b\in B\}$. Then for each $\xi\geq\alpha$ and $b\in B$ we have $bl_{\xi}=bl_{\alpha}$. Since $z\in\overline{B}$, for each open neighborhood $U$ of $z$ and $\xi\geq\alpha$ we have $Ul_{\alpha}\cap Ul_{\xi}\neq\varnothing$. Taking into account that $X$ is a Hausdorff semitopological semilattice, we conclude that $zl_{\alpha}=zl_{\xi}$ for all $\xi\geq\alpha$. It follows that $|zL|<\omega_{1}$ and, as such, $z\in I_{L}$. ∎
@@
Let $X$ be a countably tight semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\max)$. Then $I_{L}$ is a closed ideal.
Note that the order topologies on semilattices $(\omega_{1},\max)$ and $(\omega_{1},\min)$ coincide. The following result follows from [2, Theorem 1]
@@
Let $L=\{l_{\alpha}:\alpha<\omega_{1}\}\subseteq X$, where $l_{\alpha}\leq l_{\beta}$ if and only if $\alpha\leq\beta$. It is clear that $L\subset I_{L}$. Fix any $x\in I_{L}$. Since $xL$ is countable, there exists $\delta<\omega_{1}$ such that $xl_{\xi}=xl_{\delta}$ for each $\xi\geq\delta$. By Lemma 3.11, the chain $\overline{L}$ is topologically isomorphic to $(\omega_{1},\min)$ equipped with the order topology, and contains $L$ as a cofinal subset. Hence $xl_{\delta}=xa$ for each $a\in\overline{L}\cap{\uparrow}l_{\delta}$, implying that the set $x\overline{L}$ is countable. It follows that $I_{L}=I_{\overline{L}}$. Since $X$ is separable, it contains a countable dense subsemilattice $D$. Seeking a contradiction, assume that there exists an open set $U\supseteq L$ such that $U\subset I_{L}$. Consider the subsemilattice $Y$ of $X$ generated by $\overline{L}\cup(D\cap U)$. Since $U\subseteq I_{L}=I_{\overline{L}}$ we obtain that the set
@@
is countable. Since $D$ is dense in $X$ and $U\supseteq L$ is open, we get that $L\subset\overline{U\cap D}$. It follows that $\overline{L}\subset\overline{U\cap D}$. Thus the countable set $D^{\prime}$ is dense in the topological semilattice $Y$, which contradicts Theorem 4.6. ∎

### v2

@@ 2 Compact-like semigroups with dense inverse subsemigroups and the automatic continuity of inversion
$xx^{-1}=xx^{-2}x^{2}x^{-2}x^{2}x^{-1}$ $=xx^{-2}x^{2}x^{-1}=$
$=(xx^{-1})(x^{-1}x)(xx^{-1})=(xx^{-1})(x^{-1}x).$
Hence $xx^{-1}=x^{-1}x$ witnessing that the semigroup $S$ is Clifford. ∎
By $\beta X$ we denote the Stone-Čech compactification of a Tychonoff space $X$ (see [23, Chapter 3.6]). The following result appears in [4, Theorem 2.3].
@@
We are in a position to prove Theorem 1.8. We need to show that inversion is continuous in any inverse topological semigroup $S$ provided (i) $S{\times}S$ is pseudocompact, or (ii) $S$ is regular, topologically periodic and $S{\times}S$ is countably compact.
Proof of Theorem 1.8.
(i) By Theorem 2.9, $\beta S$ is a compact topological semigroup that contains $S$ as a dense inverse subsemigroup. Theorem 1.6 implies that $\beta S$ is a topological inverse semigroup. Then the continuity of inversion in $\beta S$ yields the continuity of inversion in $S$.
@@
3 Chains in topological semilattices
Recall that each semilattice possesses the natural partial order $\leq$ defined by $a\leq b$ if and only if $ab=ba=a$. Further, if a semilattice is treated as a poset, then by default it is assumed to carry the natural partial order. For an element $x$ of a semilattice $X$ put ${\downarrow}x=\{y\in X\colon y\leq x\}$ and ${\uparrow}x=\{y\in X\colon x\leq y\}$. Note that if $X$ is a semitopological semilattice, then for every $x\in X$ the sets ${\downarrow}x$ and ${\uparrow}x$ are closed. A semitopological semilattice $X$ is called chain-compact if all maximal chains in $X$ are compact. The following two theorems follow from [2, Theorem 3.1].
@@ 3 Chains in topological semilattices
Let $X$ be a semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\min)$. If $\overline{L}$ is compact, then there exists $z=\sup L\in\overline{L}$ such that for each neighborhood $U$ of $z$ the set $L\setminus U$ is countable.
@@
By Proposition 3.3, $\overline{L}$ is a compact chain. Theorem 3.1(2) applied to $\overline{L}$ implies that $\overline{L}$ contains the supremum of $L$ which we denote by $z$. It is clear that $z=\sup L^{\prime}$ for each uncountable subset $L^{\prime}$ of $L$. By Theorem 3.1(2), $z\in\overline{L^{\prime}}$ for each uncountable subset $L^{\prime}$ of $L$. It follows that for each neighborhood $U$ of $z$ the set $L\setminus U$ is countable, as required. ∎
@@
Let $X$ be a semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\max)$. If $\overline{L}$ is compact, then there exists $z=\inf L\in\overline{L}$ such that for each neighborhood $U$ of $z$ the set $L\setminus U$ is countable.
@@
Let $L$ be a chain isomorphic to $(\omega_{1},\min)$. If $\overline{L}$ is compact, then Lemma 3.4 yields $z=\sup L\in\overline{L}$. The countable tightness of $X$ implies that there exists a countable subset $A$ of $L$ such that $z\in\overline{A}$. Then there exists $l\in L$ such that $A\subseteq{\downarrow}l$. Since $X$ is a semitopological semilattice, $X\setminus{\downarrow}l$ is a neighborhood of $z$ disjoint with $A$, a contradiction. In case when $L$ is isomorphic to $(\omega_{1},\max)$, the proof is similar. ∎
We are in a position to prove Theorem 1.5. We need to show that for a Tychonoff Nyikos topological semilattice $X$ the following conditions are equivalent: (i) $X$ is compact; (ii) no chain in $X$ is isomorphic to $(\omega_{1},\min)$ or $(\omega_{1},\max)$; (iii) no chain in $X$ is topologically isomorphic to $(\omega_{1},\min)$ or $(\omega_{1},\max)$ equipped with the order topology.
Proof of Theorem 1.5.
@@
(iii) $\Rightarrow$ (i): Since $X$ is first-countable, for any cardinal $\kappa>\omega_{1}$ there is no chain in $X$ that is topologically isomorphic to $(\kappa,\min)$ or $(\kappa,\max)$ endowed with the order topology. Since $X$ is countably compact, no closed chain in $X$ is topologically isomorphic to the discrete semilattices $(\omega,\min)$ and $(\omega,\max)$. By assumption, no chain in $X$ is topologically isomorphic to $(\omega_{1},\min)$ or $(\omega_{1},\max)$ endowed with the order topology. Theorem 3.1(3) implies that the semilattice $X$ is chain-compact. By [23, Corollary 3.10.15], $X{\times}X$ is countable compact and thus pseudocompact. By Theorem 2.9, $\beta X$ is a topological semigroup that contains $X$ as a dense subsemilattice. Theorem 3.2 implies that $X$ is closed in $\beta X$ and thus $X=\beta X$ is compact. ∎
The rest of this section contains technical results on chains in topological semilattices. They will be used to prove the compactness of locally compact Nyikos topological semilattices (see Theorem 4.12), which is a milestone in the proof of Main Theorem 1.4.
@@
    For each neighborhood $U$ of $x$ the set $C\setminus U$ is finite;
@@
(2) $\Rightarrow$ (3): Assume that $x\in\overline{C}$. Fix any $c\in C$ and neighborhood $U$ of $x$. Note that the set ${\uparrow}c\cap U$ is nonempty, witnessing that $c\in cU$. Since $U$ was arbitrarily chosen and $X$ is a Hausdorff semitopological semilattice, we get that $cx=c$. Since the point $c$ was chosen arbitrarily, we obtain that $c\leq x$ for all $c\in C$, i.e. $x$ is an upper bound of $C$. Assume that $y$ is another upper bound of $C$. Taking into account that $x\in\overline{C}$, we have $U\cap yU\neq\varnothing$ for each neighborhood $U$ of $x$. Since $X$ is a Hausdorff semitopological semilattice, $x=yx$. Hence $x=\sup C$.
(3) $\Rightarrow$ (1): To derive a contradiction, assume that there exists a neighborhood $U$ of $x$ such that the set $D=C\setminus U$ is infinite. It is clear that $D$ is a cofinal subset of $C$ isomorphic to $(\omega,\min)$. In particular, we have $\sup D=x$. By the countable compactness of $X$, there exists $y\in\overline{D}$. Repeating the arguments above, we obtain $y=\sup D$. Hence $y=x$, but $x\notin\overline{D}$, a contradiction. ∎
@@
Let $X$ be a semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\alpha,\min)$ for some ordinal $\alpha$. If $\overline{L}$ does not contain maximum, then $L$ is cofinal in $\overline{L}$ and $\overline{L}$ is isomorphic to $(\alpha,\min)$.
@@
Let $L=\{l_{\xi}:\xi\in\alpha\}$, where $l_{\xi}\leq l_{\eta}$ if and only if $\xi\leq\eta$. By Proposition 3.3, $\overline{L}$ is a chain. Fix any $x\in\overline{L}$. If $L\subseteq{\downarrow}x$, then $\overline{L}\subseteq{\downarrow}x$, as ${\downarrow}x$ is closed in $X$. But then $x$ is the maximum of $\overline{L}$, a contradiction. So, for each $x\in\overline{L}$, the set $L\setminus{\downarrow}x$ is nonempty. It follows that the ordinal $\alpha$ is limit and $L$ is cofinal in $\overline{L}$. For every $x\in\overline{L}\setminus L$ let $\delta(x)=\min\{\xi\in\alpha:x\leq l_{\xi}\}$. Since $L$ is cofinal in $\overline{L}$, the ordinals $\delta(x)$, $x\in\overline{L}\setminus L$, are well defined.
Seeking a contradiction, assume that there exists an infinite decreasing chain $\{x_{n}:n\in\omega\}\subseteq\overline{L}$. The chain $L$, being isomorphic to an ordinal, contains only finite decreasing chains. Hence we lose no generality assuming that $\{x_{n}:n\in\omega\}\subseteq\overline{L}\setminus L$. Observe that $\{\delta(x_{n}):n\in\omega\}$ is a nonincreasing chain in $\alpha$. Thus there exists $k\in\omega$ such that $\delta(x_{n})=\delta(x_{m})$ for each $n,m\geq k$. Then the set $W=X\setminus({\uparrow}x_{k}\cup{\downarrow}x_{k+2})$ is a neighborhood of $x_{k+1}$ which is disjoint with $L$. The obtained contradiction implies that $\overline{L}$ contains only finite decreasing chains. Hence $\overline{L}$ is isomorphic to $(\theta,\min)$ for some ordinal $\theta$. Fix the order isomorphism $\psi:\overline{L}\rightarrow\theta$. It is easy to see that $\psi$ can be defined recursively as follows: $\psi(l_{0})=0$ and for each $y\in\overline{L}$,
$\psi(y)=\begin{cases}\psi(x)+1,&\hbox{if }\exists x<y\hbox{ such that }\{z\in\overline{L}:x<z<y\}=\varnothing;\\ \sup\{\psi(x):x<y\},&\hbox{otherwise}.\end{cases}$
@@
Let $X$ be a countably compact countably tight semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\min)$. Then $\overline{L}$ is topologically isomorphic to $(\omega_{1},\min)$ endowed with the order topology, and $L$ is cofinal in $\overline{L}$.
@@
Fix any $x\in\overline{L}\setminus L$. Since $X$ is countably tight, there exists a countable subset $A$ of $L$ such that $x\in\overline{L}$. It is clear that there exists $l\in L$ such that $A\subseteq{\downarrow}l$. Since $X$ is a semitopological semillatice, the set ${\downarrow}l$ is closed and hence $x\in\overline{A}\subseteq{\downarrow}l$. It follows that for each $x\in\overline{L}$ there exists $l\in L$ with $x\leq l$. Thus, $\overline{L}$ does not contain maximum. Lemma 3.9 implies that $\overline{L}$ is isomorphic to $(\omega_{1},\min)$ and $L$ is cofinal in $\overline{L}$. Let $\overline{L}=\{y_{\xi}:\xi<\omega_{1}\}$, where $y_{\xi}\leq y_{\delta}$ if and only if $\xi\leq\delta$. Fix any $\xi<\omega_{1}$. Clearly, for each $\delta<\xi$ the set
@@
is open in $\overline{L}$. Hence the order topology on $\overline{L}$ is contained in the original one. To derive a contradiction, assume that the order topology on $\overline{L}$ is strictly coarser than the original one. Then, taking into account that for each successor ordinal $\theta<\omega_{1}$ the point $y_{\theta}$ is isolated in both topologies, there exists a limit ordinal $\xi<\omega_{1}$ and a neighborhood $U$ of $y_{\xi}$ such that for any $\delta<\xi$ the set $\{y_{\alpha}:\delta<\alpha<\xi+1\}\setminus U$ is not empty. Since $\xi$ is countable, there exists a cofinal in ${\downarrow}y_{\xi}\setminus\{y_{\xi}\}$ subset $\{\pi_{n}:n\in\omega\}\subseteq\overline{L}\setminus U$. But this contradicts Lemma 3.8, as $y_{\xi}=\sup\{\pi_{n}:n\in\omega\}$, but $y_{\xi}\notin\overline{\{\pi_{n}:n\in\omega\}}$. ∎
4 Nyikos semilattices
@@
Let $D$ be a countable dense subset of a countably compact semitopological semilattice $X$. Recall that for each $x\in X$ the set ${\downarrow}x$ is closed. Since $X$ is countably compact and the family $\{{\downarrow}x:x\in X\}$ has the finite intersection property, $Z=\bigcap_{d\in D}{\downarrow}d\neq\varnothing$. Pick any $z\in Z$ and note that $zX=z\overline{D}\subseteq\overline{zD}=\overline{\{z\}}=\{z\}$. Hence $z$ is the minimum of $X$. ∎
@@
For every semilattice $X$ and chain $L\subseteq X$ the set $I_{L}$ is either empty or an ideal.
@@
Suppose that $I_{L}\neq\varnothing$. Fix any $x\in X$ and $y\in I_{L}$. Since $|yL|<|L|$, we get $|xyL|=|x(yL)|\leq|yL|<|L|$. Hence $xy\in I_{L}$, witnessing that $I_{L}$ is an ideal. ∎
@@
Observe that $L\subseteq I_{L}$ and hence $I_{L}\neq\varnothing$. By Lemma 4.3, $I_{L}$ is an ideal. Let $L=\{l_{\xi}:\xi<\omega_{1}\}$, where $l_{\xi}\leq l_{\beta}$ if and only if $\xi\leq\beta$. Fix any adherent point $z$ of $I_{L}$. Since $X$ is countably tight, there exists a countable subset $B\subseteq I_{L}$ such that $z\in\overline{B}$. For each $b\in B$ there exists an ordinal $\alpha_{b}<\omega_{1}$ such that $bl_{\xi}=bl_{\alpha_{b}}$ for all $\xi\geq\alpha_{b}$. Let $\alpha=\sup\{\alpha_{b}:b\in B\}$. Then for each $\xi\geq\alpha$ and $b\in B$ we have $bl_{\xi}=bl_{\alpha}$. Since $z\in\overline{B}$, for each neighborhood $U$ of $z$ and $\xi\geq\alpha$ we have $Ul_{\alpha}\cap Ul_{\xi}\neq\varnothing$. Taking into account that $X$ is a Hausdorff semitopological semilattice, we conclude that $zl_{\alpha}=zl_{\xi}$ for all $\xi\geq\alpha$. It follows that $|zL|<\omega_{1}$ and hence $z\in I_{L}$. ∎
@@
Let $X$ be a countably tight semitopological semilattice and $L\subseteq X$ be a chain isomorphic to $(\omega_{1},\max)$ such that $I_{L}\neq\varnothing$. Then $I_{L}$ is a closed ideal.
Note that the order topologies on semilattices $(\omega_{1},\max)$ and $(\omega_{1},\min)$ coincide. The following result follows from [3, Theorem 1]
@@
Let $L=\{l_{\alpha}:\alpha<\omega_{1}\}\subseteq X$, where $l_{\alpha}\leq l_{\beta}$ if and only if $\alpha\leq\beta$. It is clear that $L\subseteq I_{L}$. Fix any $x\in I_{L}$. Since $xL$ is countable, there exists $\delta<\omega_{1}$ such that $xl_{\xi}=xl_{\delta}$ for each $\xi\geq\delta$. By Lemma 3.11, the chain $\overline{L}$ is topologically isomorphic to $(\omega_{1},\min)$ equipped with the order topology, and contains $L$ as a cofinal subset. Hence $xl_{\delta}=xa$ for each $a\in\overline{L}\cap{\uparrow}l_{\delta}$, implying that the set $x\overline{L}$ is countable. It follows that $I_{L}=I_{\overline{L}}$. Since $X$ is separable, it contains a countable dense subsemilattice $D$. Seeking a contradiction, assume that there exists an open set $U\supseteq L$ such that $U\subseteq I_{L}$. Consider the subsemilattice $Y$ of $X$ generated by $\overline{L}\cup(D\cap U)$. Since $U\subseteq I_{L}=I_{\overline{L}}$ we obtain that the set
@@
is countable. Since $D$ is dense in $X$ and $U\supseteq L$ is open, we get that $L\subseteq\overline{U\cap D}$. It follows that $\overline{L}\subseteq\overline{U\cap D}$. Thus the countable set $D^{\prime}$ is dense in the topological semilattice $Y$, which contradicts Theorem 4.6. ∎